# Multi-Sensor Diffusion-Driven Optical Image Translation for Large-Scale Applications

### v3

@@ V-A1 Comparison Metrics
Looking at the numbers of Table I, we observe that the full proposed technique, Conditional DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting, reaches, by a good margin, better results in mLPIPS and FID metrics when compared to all other methods and to the Sentinel-II baseline. FID dropped by 52.44 points when compared to Sentinel-II images—an expressive improvement of $53.46\%$. When compared to Regression and Regression + $\mathbf{W}$, the FID has been respectively reduced by 18.48 and by 21.8 points—an improvement of $28.82\%$ and $31.92\%$. Among the state-of-the-art methods, only SwinIR and SRDenseNet reached comparable mLIPS and FID values to the ones from our approach.
@@
Interestingly, the vanilla DDIM-based models (Conditional DDIM and SR3) lose to all methods in any of the metrics. Even worse, they lose to the original images in both mLPIPS and mPSNR. The empirical reason for this is that the vanilla diffusion training seems to get confused by the multiple different overall tonalities of patches containing similar features, which impacts the convergence ability of the model. Deck and Bischoff [deck2023easing] noticed that diffusion models tend to present such color shifting problems, which tends to get worse for larger images. This issue seems to be eliminated when whitening and coloring procedures are used, as evidenced by the performance superiority of both Conditional DDIM + $\mathbf{W},\mathbf{C}$ and Conditional DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting. Looking at mLPIPS results, we observe that the regression-based models beat by a small margin the original Sentinel-II images, whereas Conditional DDIM + $\mathbf{W},\mathbf{C}$ and Conditional DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting show a much more significant improvement in the same regard. This indicates that, in a patch-by-patch basis, only DDIM + $\mathbf{W},\mathbf{C}$ and DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting models were able to heavily reduce the perceptual distance of patches when compared to their Planet Dove versions. The proposed method with PSNR Voting has reduced mLPIPS by 0.1093 points, meaning that the produced patches are, on average, $36.71\%$ more perceptually similar to their corresponding Planet Dove patches than their Sentinel-II versions. When we compare the last two rows of the table, PSNR voting brings a significant improvement in all metrics: $-5.469\%$ for mLPIPS, $-2.665\%$ for FID, and $+2.144\%$ for mPSNR. We observed, moreover, that PSNR voting is specially useful to discard rare but existent undesired model hallucinations and poor translations, which would negatively impact, for example, change detection results. These improvements, due to the increased number of model evaluations, come at the cost of higher numerical complexity.
@@
While regression models achieve higher mPSNR values, as discussed, PSNR’s preference for blurry regression-produced outputs does not align well with human perception. The same can be said about the SwinIR, ShuffleMixer and SRDenseNet methods, as they are also trained using regression objectives. Consequently, learning-based metrics, like the FID and the LPIPS, lately have been the standard in image generation comparisons, rendering the PSNR performance superiority of regression models irrelevant.
As discussed, the very poor performance shown by the SR3 model is in part due to what plagued the performance of the classically trained DDIM model. However, even compared to it, the SR3 model presented far worse numbers. This fact shows how the model architecture chosen for our method has been able to better extract the complex features of the conditioning input and generate better translations.
@@ VI Discussion
The results of our study demonstrate the efficacy of leveraging Denoising Diffusion Implicit Models (DDIM) for large-scale optical image translation. Our approach successfully addresses the limitations of existing Image-to-Image (I2I) methods, particularly in handling large-scale multi-patch imagery. By focusing on super-resolving low spatial resolution images into high-resolution equivalents, our method achieves significant improvements in both radiometric accuracy and feature representation.
@@
While our method effectively handles large-scale imagery, the computational complexity remains challenging. Future work will also explore optimization techniques to reduce the computational load without compromising the model’s performance. Additionally, extending our approach to other types of remote sensing data, such as radar or hyperspectral images, could broaden its applicability and impact. Another potential direction for future work is integrating our method with real-time processing systems. This would enable the rapid analysis of remote sensing data, particularly valuable in time-sensitive applications such as disaster response and environmental monitoring. Furthermore, incorporating advanced machine learning techniques, such as reinforcement learning, could enhance the adaptability and robustness of our method in diverse and dynamic environments.
In conclusion, this paper presented a novel and effective DDM-based approach for large-scale optical image translation addressing key challenges. The results demonstrate significant improvements in image quality and consistency, paving the way for more accurate and reliable remote sensing applications.
@@ Appendix A Implementation Details
The model chosen for the experiments is the one presented in [Wang2022], where the authors propose a U-Net-style model that uses attention mechanisms to improve image fidelity. Differently from the diffusion-based models presented in [Saharia2021, Saharia2021SR3, ho2022classifierfree], this model processes the noisy input $\bm{y}_{t}$ and the conditioning input $\bm{x}$ independently. That is $\bm{y}_{t}$ is fed right into the encoder and is processed normally, whereas $\bm{x}$ is re-introduced into the decoder by inserting it into each residual block. To properly match the spatial resolution required by each residual block, $\bm{x}$ is processed by spatially adaptive normalization (SPADE) blocks. As confirmed by the performed experiments, we observed that this modification improved significantly the generation quality and fidelity to the conditioning input. Some internal parameters of the model were adjusted to maximize performance while limiting its size due to hardware constraints. The number of channels has been set to be a multiple of $96$. The number of residual blocks for each downsampling operation has been set to $2$. Attention resolutions were set to $2,4,8$ and $16$. Finally, the number of attention heads has been set to $4$. This configuration resulted in a model with around 195 million trainable parameters. For training, batch size has been set to $5$, and the learning rate has been set to $10^{-4}$.

### v4

@@ V-A1 Comparison Metrics
Looking at the numbers of Table I, we observe that the full proposed technique, Conditional DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting, reaches, by a good margin, better results in mLPIPS and FID metrics when compared to all other methods and to the Sentinel-II baseline. FID dropped by 52.44 points when compared to Sentinel-II images—an expressive improvement of $53.46\%$. When compared to Regression and Regression + $\mathbf{W}$, the FID has been respectively reduced by 18.48 and by 21.8 points—an improvement of $28.82\%$ and $31.92\%$. Among the state-of-the-art methods, only SwinIR and SRDenseNet reached comparable mLPIPS and FID values to the ones from our approach.
Interestingly, the vanilla conditional DDIM-based tested models—using SDM and SR3 backbones——perform worse than all other methods in all metrics. They even underperform compared to the original images in both mLPIPS and mPSNR. The empirical reason is that the vanilla diffusion training seems confused by the multiple different overall tonalities of patches containing similar features, impacting the model’s convergence ability. Deck and Bischoff [deck2023easing] noticed that diffusion models tend to present such color-shifting problems, which worsen for larger images.
@@
This issue is mitigated when whitening and coloring procedures are used, as evidenced by the performance superiority of both Conditional DDIM + $\mathbf{W},\mathbf{C}$ and Conditional DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting (Proposed). Looking at mLPIPS results, we observe that the regression-based models slightly outperform the original Sentinel-II images, whereas Conditional DDIM + $\mathbf{W},\mathbf{C}$ and Conditional DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting show a much more significant improvement. This indicates that, in a patch-by-patch basis, only DDIM + $\mathbf{W},\mathbf{C}$ and DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting models were able to heavily reduce the perceptual distance of patches when compared to their Planet Dove versions. The proposed method with PSNR Voting has reduced mLPIPS by 0.1093 points, meaning that the produced patches are, on average, $36.71\%$ more perceptually similar to their corresponding Planet Dove patches than their Sentinel-II versions. Comparing the last two rows of the table, PSNR voting brings a significant improvement in all metrics: $-5.469\%$ for mLPIPS, $-2.665\%$ for FID, and $+2.144\%$ for mPSNR. Moreover, we observed that PSNR voting is specially useful to discard rare but existent undesired model hallucinations and poor translations, which would negatively impact, for example, change detection results. These improvements, due to the increased number of model evaluations, come at the cost of higher numerical complexity.
@@
While regression models achieve higher mPSNR values, PSNR’s preference for blurry regression-produced outputs does not align well with human perception. The same applies to the SwinIR, ShuffleMixer, and SRDenseNet methods, as they were also trained using regression objectives. Consequently, learning-based metrics like FID and LPIPS have become the standard in image generation comparisons, rendering the PSNR performance superiority of regression models less relevant.
As discussed, the very poor performance shown by the conditional DDIM using the SR3 backbone is in part due to what plagued the performance of the classically trained DDIM model using the SDM backbone. However, even compared to it, the SR3 model presented far worse numbers. This fact shows how the model architecture chosen for our method has been able to better extract the complex features of the conditioning input and generate better translations. To further explore the influence of the backbone architecture, we also tested the proposed method using SR3 as the backbone, applying our forward and reverse diffusion methods. This dramatically improved the performance compared to its vanilla DDIM training, indicating that our approach enhances the capabilities across different network architectures. However, when compared to using the SDM backbone, the proposed method with the SR3 backbone performed much worse in both mLPIPS and FID, and could not effectively super-resolve the images. This suggests that while our proposed method improves performance for different backbones, the choice of architecture remains critical for achieving optimal results. The SDM backbone has been better able to extract complex features from the conditioning input and generate superior translations, highlighting the importance of selecting an appropriate architecture in conjunction with our proposed methods. Further details about the SDM backbone and underlying hyperparameters are discussed in Appendix A.
@@ VI Discussion
The results of our study demonstrate the efficacy of leveraging our proposed Denoising Diffusion method for large-scale optical image translation. Our approach successfully addresses the limitations of existing Image-to-Image (I2I) methods, particularly in handling large-scale multi-patch imagery. By focusing on super-resolving low spatial resolution images into high-resolution equivalents, our method achieves significant improvements in both radiometric accuracy and feature representation.
@@
As described in Section III, when a domain B HR image is not available for color extraction during inference, it is possible to extract the needed radiometry information directly from the input LR domain A to restore the overall colorization of the domain B output. Given the overall radiometric shifts between these domains, relying on overall radiometry information from domain A may result in synthetic images that may not perfectly represent the radiometry expected from the target domain. However, in cases where radiometry information from domain B is not available, this compromise can still produce images that are adequate for most downstream tasks, seen that the radiometric shifts between domains are rarely large enough to cause problems.
While our method effectively handles large-scale imagery, the computational complexity remains challenging. Future work will also explore optimization techniques to reduce the computational load without compromising the model’s performance. Additionally, extending our approach to other types of remote sensing data, such as radar or hyperspectral images, could broaden its applicability and impact. Notably, when the number of channels ($n_{\text{ch}}$) is reduced to one, as in panchromatic or single-band SAR images, the whitening and colorization processes become redundant due to the subsequent normalization. In such cases, an alternative approach would be to remove the $[-1,1]$ normalization and instead normalize the values of all patches by dividing them by a fixed maximum value to keep them within a reasonable range. This would be followed by mean subtraction and posterior mean addition in the resulting outputs. We did not, however, performed these tests, as working with panchromatic images is out of the scope of the manuscript.
@@ VII Conclusion
We introduced an innovative deep learning-based method that employs denoising diffusion models to generate synthetic high-spatial-resolution satellite images from low-spatial-resolution inputs of a different sensor. Our method incorporates novel forward and reverse diffusion procedures, including color standardization and initial noise selection techniques, which align input and output patches and reduce model hallucinations and color inconsistencies. Trained and tested on a large and diverse dataset of paired images, our method demonstrated superior performance compared to existing solutions, including the popular classifier-free guided DDIM framework.
We also applied our method to the task of heterogeneous change detection between images from different sensors, demonstrating its effectiveness in two urban areas: Beirut, Lebanon, and Austin, USA. While our method effectively handles large-scale imagery and maintains uniformity across patches, certain limitations remain—such as when color information from the target domain is unavailable during inference, and the high computational inference complexity. To enhance the robustness of our approach, future work will focus on addressing these limitations. We plan to extend our method to other combinations of sensors and investigate the use of synthetic images in target change detection.
@@ Appendix A Implementation Details
SDM [Wang2022], the chosen backbone for the proposed method’s experiments, is a U-Net-style model that uses attention mechanisms to improve image fidelity. Differently from the diffusion-based models presented in [Saharia2021, Saharia2021SR3, ho2022classifierfree], this model processes the noisy input $\bm{y}_{t}$ and the conditioning input $\bm{x}$ independently. That is $\bm{y}_{t}$ is fed right into the encoder and is processed normally, whereas $\bm{x}$ is re-introduced into the decoder by inserting it into each residual block. To properly match the spatial resolution required by each residual block, $\bm{x}$ is processed by spatially adaptive normalization (SPADE) blocks. As confirmed by the performed experiments, we observed that this modification improved significantly the generation quality and fidelity to the conditioning input. Some internal parameters of the model were adjusted to maximize performance while limiting its size due to hardware constraints. The number of channels has been set to be a multiple of $96$. The number of residual blocks for each downsampling operation has been set to $2$. Attention resolutions were set to $2,4,8$ and $16$. Finally, the number of attention heads has been set to $4$. This configuration resulted in a model with around 195 million trainable parameters. For training, batch size has been set to $5$, and the learning rate has been set to $10^{-4}$.
@@ B-E Min-SNR Weighting Strategy
[Uncaptioned image] João Gabriel Vinholi received his Engineering degree in Electrical Engineering from the Federal University of Santa Catarina (UFSC), Florianópolis, Brazil, in 2020. From 2021 to 2023, he worked at ICEYE, Luxembourg, as a machine learning engineer. During his tenure, he focused on developing deep learning-based algorithms for multi-temporal flood mapping, deforestation mapping, building damage detection, and anomaly detection in X-band SAR imagery. He also specialized in MLOps for deploying these algorithms efficiently. Since 2023, he has been with the Luxembourg Institute of Science and Technology (LIST), Esch-sur-Alzette, Luxembourg, as a machine learning scientist, where he continues to conduct research and development of machine learning-based tools for remote sensing applications. Concurrently, since 2021, Vinholi has been a Doctoral Candidate of Electronics and Computer Engineering at the Aeronautics Institute of Technology (ITA), São José dos Campos, Brazil. His research interests lie in the fields of machine and deep learning, computer vision, digital signal processing, and their applications in remote sensing.
[Uncaptioned image] Marco Chini (Senior Member, IEEE) received the M.S. degree in electronic engineering from the Sapienza University of Rome, Rome, Italy, in 2003, and the Ph.D. degree in geophysics from the University of Bologna, Bologna, Italy, in 2008. From 2003 to 2012, he collaborated with the Department of Information Engineering, Electronics and Telecommunications, Sapienza University of Rome. From 2006 and 2008, he was a Visiting Researcher with the Department of Aerospace Engineering Science, University of Colorado, Boulder, CO, USA, where he was involved in very high-resolution optical images for urban monitoring. From 2008 to 2012, he was with the Remote Sensing Group, Istituto Nazionale di Geofisica e Vulcanologia, Rome, where his research topic was the development of Earth Observation (EO)-based innovative algorithms for detecting damages caused by earthquakes. Since 2013, he has been with the Luxembourg Institute of Science and Technology (LIST), Esch-sur-Alzette, Luxembourg, formerly the Centre de Recherche Public—Gabriel Lippmann. He has been involved in projects for mapping flooded areas, detecting damage caused by earthquakes, monitoring urban sprawl, detecting surface changes, and identifying risk-/hazard-prone areas. His research interests include analysis of multitemporal data, deep learning, feature extraction, data fusion and segmentation using synthetic aperture radar and optical data, soil moisture retrieval, and SAR interferometry technique for geophysical applications. Dr. Chini is an Associate Editor of the IEEE TRANSACTIONS ON GEOSCIENCE AND REMOTE SENSING, an Editorial Advisor of Springer’s publication program in the Remote Sensing/Photogrammetry, and an Associate Editor of the European Journal of Remote Sensing.
[Uncaptioned image] Anis Amziane received the M. S. degree in Signal and Image processing from the University of Burgundy, Dijon, France, in 2018, and the Ph.D. degree in Computer Engineering and Signal/Image processing from the University of Lille, Lille, France, in 2022. From 2021 to 2022, he was a Research and Teaching Associate at the Faculty of Science and Technology of the University of Lille, France. In 2023, he joined the Luxembourg Institute of Science and Technology (LIST), Esch-sur-Alzette, Luxembourg, as a Postdoctoral researcher, where he conducts the research and development of machine learning algorithms for remote sensing image super resolution. His research interests include multispectral proximal/remote imaging, illumination-robust feature extraction, spectral filter array (SFA) image analysis, and image super resolution.
[Uncaptioned image] Renato Machado received the B.S.E.E. degree from the São Paulo State University (UNESP), Ilha Solteira, SP, Brazil, in 2001. He received the M.Sc. and Ph.D. degrees in Electrical Engineering from the Federal University of Santa Catarina (UFSC), Florianópolis, SC, Brazil, in 2004 and 2008, respectively. From November 2013 to February 2015, Dr. Machado was a visiting Research Fellow at Blekinge Institute of Technology (BTH) in partnership with Saab AB. From August 2009 to December 2017, he was an Assistant (2008-2016) and Associate (2017) Professor at the Federal University of Santa Maria, RS, Brazil, where he lectured many courses in bachelor and graduate programs and assumed different positions in the institution, namely, Researcher leader of the Communications and Signal Processing Research Group, Coordinator of the Telecommunications Engineering Program, and Director of the Aerospace Science Laboratory. Since December 2017, Dr. Machado has been an Associate Professor at the Aeronautics Institute of Technology (ITA). He is the Research leader of the Digital and Signal Processing Laboratory, ITA, and SAR and Radar Signal Processing Laboratory, ITA. Prof. Machado is a Principal Investigator of BI0S (Brazilian Institute of Data Science), a Research Center for Applied Science and Technology in Artificial Intelligence. - FAPESP/MCTI/CGI. Between August 2021 and July 2022, he was the head of the Graduate Program in Electronics and Computer Engineering, ITA. Since June 2022, Prof. Machado has been the head of the Electronic Engineering Division, ITA. In February 2021, Dr. Machado received the Senior Membership degree from the Institute of Electrical and Electronics Engineers (IEEE). His research interests include SAR Processing, SAR Image Processing, Change Detection, Radar Signal Processing, and Digital Signal Processing.
[Uncaptioned image] Danilo Silva received the B.Sc. degree from the Federal University of Pernambuco (UFPE), Recife, Brazil, in 2002, the M.Sc. degree from the Pontifical Catholic University of Rio de Janeiro (PUC-Rio), Rio de Janeiro, Brazil, in 2005, and the Ph.D. degree from the University of Toronto, Toronto, Canada, in 2009, all in electrical engineering. From 2009 to 2010, he was a Postdoctoral Fellow at the University of Toronto, at the École Polytechnique Fédérale de Lausanne (EPFL), and at the State University of Campinas (UNICAMP). In 2010, he joined the Department of Electrical and Electronic Engineering, Federal University of Santa Catarina (UFSC), Florianópolis, Brazil, where he is currently an Associate Professor. His current research interests include machine learning, computer vision, signal processing, and their applications. His previous research interests include wireless communications, network coding, channel coding, and information theory. Dr. Silva is a member of the Brazilian Telecommunications Society (SBrT). He was a recipient of a CAPES Ph.D. Scholarship in 2005, the Shahid U.H. Qureshi Memorial Scholarship in 2009, and a FAPESP Postdoctoral Scholarship in 2010. He also served as Technical Program Co-Chair of the XXXVIII Brazilian Symposium on Telecommunications and Signal Processing (SBrT 2020).
[Uncaptioned image] Patrick Matgen Patrick Matgen earned his M.Sc. in Environmental Engineering at the Ecole Polytechnique Fédérale de Lausanne (Switzerland) in 2002. In 2011 he finalized his Ph.D dissertation “Surface and subsurface water from space: on the integration of microwave remote sensing observations with flood prediction systems” at the Technical University in Delft (The Netherlands). As a Project Manager from 2002 and then as a Lead Research and Technology Associate from 2015 until 2020 he was responsible for acquiring funding and managing Research Development and Innovation projects in Earth Observation and hydrological-hydraulic modelling within the research unit ‘Environmental sensing and modelling’ of the Luxembourg Institute of Science and Technology (LIST). Since January 2020 he is leading the group “remote sensing and natural resources modelling” of the research unit. The group’s aim is to develop the synergistic use, processing and interpretation of data from multiple complementary active and passive sensors installed on both space- and airborne platforms and to provide deeper insights into the relations between spectral features and properties of Earth’s natural resources. Patrick Matgen’s personal research interests include the assimilation of remote sensing-derived observations into hydrodynamic models and the assessment of flood hazard and risk at large scale.